OpenStudy (shelby1290):

Describe the steps needed to graph each of the following functions: a) f(x)=-x^2 + 34 b) f(x) =-5.5(x-6)^2+11 c) f(x)=1/10x^2+16 d) f(x) = -(x+18)^2+3

OpenStudy (hwyl):

they're all quadratic, correct?

OpenStudy (shelby1290):

The steps are supposed to be described in words, i believe.

OpenStudy (hwyl):

so what do you know about quadratic functions?

OpenStudy (hwyl):

dot dot dot

OpenStudy (shelby1290):

@hwyl so what do i write down?

OpenStudy (hwyl):

read the contents of that website

OpenStudy (danjs):

depends on what kind of resolution or general shape you need

OpenStudy (danjs):

they are all quadratic, you can put them in y = ax^2 + bx + c form

OpenStudy (danjs):

most important values to graph, Vertex x-intercept if any y-intercepts if any

OpenStudy (shelby1290):

@DanJS , @hwyl i don't think i can say the same answer for all of them though

OpenStudy (danjs):

start with one of them, and explain and show how to get a basic graph

OpenStudy (danjs):

f(x) = -x^2 + 34 vertex--- x = -b/(2*a) = -0/-2 = 0 y=f(0) = 34 Vertex is at (x,y) = (0,34)

OpenStudy (shelby1290):

okay so like for example with part a) Find the vertex, create a table of values, and graph them?

OpenStudy (danjs):

the sin on the coefficient of the squarred term is negative, and tells you the parabola opens down...there will be x-intercepts x-intercepts are where y=0,, solve y = 0 = -x^2 + 34 plot the points (x,0)

OpenStudy (shelby1290):

It's just that i feel like the same steps are followed for the every question. So, i don't understand how you're supposed to find the steps for each one to differentiate in words

OpenStudy (danjs):

yeah, i would find the vertex first, then find the x and y axis intercepts if they exist

OpenStudy (danjs):

you can use numbers to, just explain what you are doing, and show the graph of each point you find

OpenStudy (danjs):

You can list the values in a table if you want, but not really needed

OpenStudy (danjs):